OpenStudy (anonymous):

Find a degree 5 Polynomial p(x) with Zeros: 1,2,3,5+i and p(6)=240

OpenStudy (anonymous):

need help with problem too?

OpenStudy (anonymous):

If one is a zero, then (x - 1) is a factor.

OpenStudy (imstuck):

If those are the zeros, then the factors from which they came are (x - 1)(x - 2)(x - 3)(x - (5 + i))

OpenStudy (imstuck):

The last one I a working on.

OpenStudy (anonymous):

ohh okay so then it would be (x-1)(x-2)(x-3)(x-5-i)?

OpenStudy (anonymous):

and i would use the synthetic division for P(6)

OpenStudy (jdoe0001):

hmmm

OpenStudy (imstuck):

P(6) means that when you put 6 in for x, your polynomial = 240. So the last factor I would say is (x - 6). Let's do that math now with all our factors. We will save the "i" one for last. Ok?

OpenStudy (jdoe0001):

keep in mind that "complex roots" do not come all by themselves they have a companion that comes with them, their "conjugate" thus x = 5+i => x -5 -i=0 has its companion of \(\bf x-5+i\) and thus would be the 5 roots

OpenStudy (jdoe0001):

or I should put them as (x-5) -i and (x-5) + i

OpenStudy (anonymous):

@IMStuck OKAY

OpenStudy (anonymous):

@jdoe0001 SO then i would have those two as well

OpenStudy (jdoe0001):

\(\bf [(x-5) -i][(x-5) + i]\qquad recall\implies {\color{brown}{ (a-b)(a+b) = a^2-b^2}} \\ \quad \\\ [(x-5) -i][(x-5) + i]\implies [(x-5)^2-i^2] \\ \quad \\ {\color{brown}{ i^2\to \sqrt{-1}\cdot \sqrt{-1}\to \sqrt{(-1)^2}\to -1}} \\ \quad \\\ [(x-5)^2-i^2]\implies [(x-5)^2-({\color{brown}{ -1}})]\implies [(x-5)^2+1]\)

OpenStudy (anonymous):

@jdoe0001 oh okay

OpenStudy (jdoe0001):

and of course, to get the polynomial, as IMStuck said. just multiply them all

OpenStudy (anonymous):

what do you mean multiply them all?

OpenStudy (jdoe0001):

\(\bf (x-1)(x-2)(x-3)[(x-5)^2+1]=0 \\ \quad \\ (x-1)(x-2)(x-3)[(x-5)^2+1]=\textit{original polynomial}\)

OpenStudy (anonymous):

okay then i would get :1,2,3,

OpenStudy (anonymous):

and (x^2-10x+26)

OpenStudy (jdoe0001):

yeap, and to get the original polynomial. you'd just multiply all roots so \(\bf (x-1)(x-2)(x-3)(x^2-10x+26)=\textit{original polynomial}\)

OpenStudy (anonymous):

SO it's x^5-16x^4+97x^3-272x^2+346x-156

OpenStudy (jdoe0001):

lemme check

OpenStudy (anonymous):

okay

OpenStudy (jdoe0001):

\(\bf {\color{brown}{ (x-1)(x-2)(x-3)}}(x^2-10x+26) \\ \quad \\ {\color{brown}{ (x^3-6x^2+11x-6)}}(x^2-10x+26)\\ x^5-6x^4+11x^3-6x^2-10x^4+60x^3-110x^2+60x\lrcorner\\ +26x^3-156x^2+286-156 \\ \quad \\ {\color{blue}{ x^5-16x^4+97x^3-272x^2+130}}\)

OpenStudy (jdoe0001):

hmmm wait.. a sec.. I'm mising the "x" bit =)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

:)

OpenStudy (jdoe0001):

hmm anyow... yes,.. your correct :)

OpenStudy (jdoe0001):

sorta ate the "x" for a sec =)

OpenStudy (jdoe0001):

\(\bf {\color{brown}{ (x-1)(x-2)(x-3)}}(x^2-10x+26) \\ \quad \\ {\color{brown}{ (x^3-6x^2+11x-6)}}(x^2-10x+26)\\ x^5-6x^4+11x^3-6x^2-10x^4+60x^3-110x^2+60x\rgroup\\ +26x^3-156x^2+286x-156 \\ \quad \\ {\color{blue}{ x^5-16x^4+97x^3-272x^2+346x-156}}\)

OpenStudy (anonymous):

so what do i do with p(6)=240

OpenStudy (jdoe0001):

hmm

OpenStudy (anonymous):

@IMStuck Hey do you know what do i do with p(6)=240

OpenStudy (jdoe0001):

so I just ran it through, setting x = 6 and gave me p(x) or p(6) \(\bf p(x)={\color{blue}{ x^5-16x^4+97x^3-272x^2+346x-156}} \\ \quad \\ p(6)=120\qquad thus \\ \quad \\ p(x)=2({\color{blue}{ x^5-16x^4+97x^3-272x^2+346x-156}}) \\ \quad \\ p(6)=2(120)\to 240\)

OpenStudy (jdoe0001):

ahemm it gave 120 obviously =)

OpenStudy (jdoe0001):

so that means the polynomial originally has a multiplier of 2 in front

OpenStudy (anonymous):

so you just found a number that gave you 240?

OpenStudy (anonymous):

why did you use those two number why not 3(80)

OpenStudy (jdoe0001):

ohh because if you plug "6" as the "input" for "x" that is, set x = 6 the polynomial of \(\bf x^5-16x^4+97x^3-272x^2+346x-156\) gives 120 but we know that p(6) is twice that much that means, that the polynomial is being multiplied by a digit of 2

OpenStudy (anonymous):

ohh okay i get it now :)

OpenStudy (anonymous):

Thanks!!

OpenStudy (jdoe0001):

yw